the perimeter is 36 inches and the length is 3 times the width. what is the length and width?
Answer by checkley77(12844)

L=3W
2L+2W=36
2(3W)+2W=36
6W+2W=36
8W=36
W=36/8
W=4.5 INCHES FOR THE WIDTH.
L=3*4.5=13.5 INCHES FOR THE LENGTH.
PROOF:
2*13.5+2*4.5=36
27+9=36
36=36